# Multi-objective Reinforcement Learning with Nonlinear Preferences: Provable Approximation for Maximizing Expected Scalarized Return

### v3

@@ Appendix A Proofs of Technical Lemmas
Lemma 4.
Let $M$ be any MOMDP, and let $\pi$ be any policy in $M$. Assume the welfare function $W:\mathbb{R}^{d}\to\mathbb{R}$ is uniformly continuous and $\mathbf{R}(s,a)\in[0,1]^{d}~{}\forall s,a$. Then for all $\epsilon>0$, there exists $\delta_{\epsilon}>0$ such that if
$T\geq\left(\frac{1}{1-\gamma}\right)\log\left(\frac{1}{\delta_{\epsilon}(1-\gamma)}\right)$
then for any state $s,$
Proof.
[…]
$=$ $\bigg{\|}\sum_{k=T+1}^{\infty}\gamma^{k-1}\mathbf{R}(s_{k},a_{k})\bigg{\|}$
$\leq$ $\gamma^{T}\left(\frac{1}{1-\gamma}\right)$
Solving
$\gamma^{T}\left(\frac{1}{1-\gamma}\right)\leq\delta_{\epsilon}$
@@
where the sum is taken over all length-$t$ trajectories $\tau^{\prime}$ beginning at $s$, and $Pr^{\pi}(\tau^{\prime})$ is the probability that $\tau^{\prime}$ is traversed under policy $\pi$. Note that we can further decompose $\tau^{\prime}$ into $\tau^{\prime}=(s,a)\oplus\tau^{\prime\prime}$ where $(s,a)$ denotes the state-action pair corresponding to the current timestep, $\oplus$ concatenates, and $\tau^{\prime\prime}$ denotes a trajectory of length $t-1$ beginning at some state $s^{\prime}$ transitioned from state $s$ via taking action $a$. Then we can rewrite $V^{*}$ as
$V^{*}(s,\tau,t)=\max_{\pi}\sum_{a}\pi(a\mid s,\tau,t)~{}\mathbb{E}_{s^{\prime}}\bigg{[}\sum_{\tau^{\prime\prime}}Pr^{\pi}(\tau^{\prime\prime})W(\mathbf{R}(\tau)+\gamma^{T-t}\mathbf{R}(s,a)+\gamma^{T-t+1}\mathbf{R}(\tau^{\prime\prime}))\bigg{]}$
@@
$V^{*}(s,\tau,t)$ $=\max_{a}\mathbb{E}_{s^{\prime}}\bigg{[}\max_{\pi}\sum_{\tau^{\prime\prime}}Pr^{\pi}(\tau^{\prime\prime})W(\mathbf{R}(\tau)+\gamma^{T-t}\mathbf{R}(s,a)+\gamma^{T-(t-1)}\mathbf{R}(\tau^{\prime\prime}))\bigg{]}$
@@
$=$ $\sum_{\tau}Pr^{\pi_{1}}[\tau]\bigg{|}W(\mathbf{R_{1}}+\mathbf{R}(\tau))-W(\mathbf{R_{2}}+\mathbf{R}(\tau))\bigg{|}$
@@
$\forall s\in\mathcal{S},\mathbf{R_{acc}}\in\mathbb{R}^{d},t\in\{0,1,\dots,T\}$, where $V(s,\mathbf{R_{acc}},t)$ is computed by Algorithm 1.
@@
We prove this additive error bound by induction on the number of timesteps $t$ remaining in the task. Cleary $V(s,\mathbf{R_{acc}},0)=W(\mathbf{R_{acc}}),\forall s$ and $\mathbf{R_{acc}}$. Suppose that $V(s,\mathbf{R_{acc}},t)\geq V^{*}(s,\mathbf{R_{acc}},t)-t\epsilon$ for all $t=0,\dots,k-1$. Then consider with $k$ steps remaining.
@@
$\geq$ $\max_{a}\sum_{s^{\prime}}Pr(s^{\prime}|s,a)\times\bigg{(}V^{*}(s^{\prime},f_{\alpha}(\mathbf{R_{acc}}+\gamma^{T-k}\mathbf{R}(s,a)),k-1)-(k-1)\epsilon\bigg{)}.$
@@
$\bigg{|}V^{*}\bigg{(}s^{\prime},f_{\alpha}(\mathbf{R_{acc}}+\gamma^{T-k}\mathbf{R}(s,a)),k-1\bigg{)}-V^{*}\bigg{(}s^{\prime},\mathbf{R_{acc}}+\gamma^{T-k}\mathbf{R}(s,a),k-1\bigg{)}\bigg{|}<\epsilon$
@@
Hence, it suffices to choose $\alpha=\delta_{\epsilon}/d$, which implies
@@
$V(s,\mathbf{R_{acc}},k)\geq$ $\max_{a}\sum_{s^{\prime}}Pr(s^{\prime}|s,a)V^{*}\bigg{(}s^{\prime},\mathbf{R_{acc}}+\gamma^{T-k}\mathbf{R}(s,a),k-1\bigg{)}-\epsilon-(k-1)\epsilon$
@@
Appendix B Ablation Studies on Accumulated Rewards Discretization
@@
We conduct further studies on the empirical effects of the discretization factor $\alpha$. For $W_{\text{SPF}}$, we set $\lambda=1$ (the smoothing factor). We gradually increase the disctretization interval of accumulated reward $\alpha$ gradually from $0.4$ to $1.2$ for a reward discount $\gamma=0.999$ for Taxi and $\gamma=0.99$ for Scavenger Hunt with the same settings as the experiments. As shown in Figure 4(a)(a) and Figure 4(b), we observe that the solution quality is often very high (better than baselines considered in the main body) even for discretization values that are substantially greater than those needed for theoretical guarantees (with $\epsilon=1$, we need $\alpha=2.5\times 10^{-3}$). Nonetheless the results are not necessarily monotone with respect to $\alpha$ in this regime.
@@
[…]
Figure 4: Ablation study on discretization factor $\alpha$.
@@
This set of ablation studies demonstrates that the RAVI algorithm is capable of finding high-quality policies for optimizing expected scalarized return despite using a larger discretization parameter than that needed for our theoretical analysis. The practical takeaway is that, if one cannot afford the runtime for a very fine $\alpha$ with worst-case theoretical guarantees, one should instead try several much larger $\alpha$ values but still less than the smoothness of the objective. Due to the superlinear runtime scaling with respect to $\alpha$ for $d>1$, this can be substantially faster than a single run with much smaller $\alpha$.
Appendix C Details on the RAEE Algorithm

### v4

@@ Appendix A Proofs of Technical Lemmas
A.1. Proof of Lemma 6
@@
where the sum is taken over all length-$t$ trajectories $\tau^{\prime}$ beginning at $s$, and $Pr^{\pi}(\tau^{\prime})$ is the probability that $\tau^{\prime}$ is traversed under policy $\pi$. Note that we can further decompose $\tau^{\prime}$ into $\tau^{\prime}=(s,a)\oplus\tau^{\prime\prime}$ where $(s,a)$ denotes the state-action pair corresponding to the current timestep, $\oplus$ concatenates, and $\tau^{\prime\prime}$ denotes a trajectory of length $t-1$ beginning at some state $s^{\prime}$ transitioned from state $s$ via taking action $a$. Since the transition to $\tau^{\prime\prime}$ is independent of earlier states/actions once the current state $s$ and action $a$ are fixed, we can simplify the optimization by shifting the summation from trajectories conditioned on $\pi$ to actions and successor states and thus rewrite $V^{*}$ as
$V^{*}(s,\tau,t)=\max_{\pi}\sum_{a}$ $\pi(a\mid s,\tau,t)$
$\mathbb{E}_{s^{\prime}}\Bigg{[}\sum_{\tau^{\prime\prime}}Pr^{\pi}(\tau^{\prime\prime})$ $W\big{(}\mathbf{R}(\tau)+\gamma^{T-t}\mathbf{R}(s,a)+\gamma^{T-t+1}\mathbf{R}(\tau^{\prime\prime})\big{)}\Bigg{]}$
@@
$V^{*}(s,\tau,t)$ $=\max_{a}\mathbb{E}_{s^{\prime}}\bigg{[}\max_{\pi}\sum_{\tau^{\prime\prime}}Pr^{\pi}(\tau^{\prime\prime})W(\mathbf{R}(\tau)$
$\qquad\qquad+\gamma^{T-t}\mathbf{R}(s,a)+\gamma^{T-(t-1)}\mathbf{R}(\tau^{\prime\prime}))\bigg{]}$
@@ A.2. Proof of Lemma 9
$\leq$ $\sum_{\tau}Pr^{\pi_{1}}[\tau]\bigg{|}W(\mathbf{R_{1}}+\mathbf{R}(\tau))-W(\mathbf{R_{2}}+\mathbf{R}(\tau))\bigg{|}$
@@ A.3. Proof of Lemma 10
$\forall s\in\mathcal{S},\mathbf{R_{acc}}\in\mathbb{R}^{d},t\in\{0,1,\dots,T\}$, where $V(s,\mathbf{R_{acc}},t)$ is computed by Algorithm 1 using $\alpha_{\epsilon}$.
@@
We prove this additive error bound by induction on the number of timesteps $t$ remaining in the task. Clearly $V(s,\mathbf{R_{acc}},0)=W(\mathbf{R_{acc}}),\forall s$ and $\mathbf{R_{acc}}$. Suppose that $V(s,\mathbf{R_{acc}},t)\geq V^{*}(s,\mathbf{R_{acc}},t)-t\epsilon$ for all $t=0,\dots,k-1$. Then consider with $k$ steps remaining.
@@
$\geq$ $\max_{a}\sum_{s^{\prime}}Pr(s^{\prime}|s,a)$
$\qquad\bigg{(}V^{*}(s^{\prime},f_{\alpha}(\mathbf{R_{acc}}+\gamma^{T-k}\mathbf{R}(s,a)),k-1)-(k-1)\epsilon\bigg{)}.$
@@
$\bigg{|}V^{*}\bigg{(}s^{\prime},f_{\alpha}(\mathbf{R_{acc}}+\gamma^{T-k}$ $\mathbf{R}(s,a)),k-1\bigg{)}$
$-V^{*}\bigg{(}s^{\prime},\mathbf{R_{acc}}+\gamma^{T-k}\mathbf{R}(s,a),k-1\bigg{)}\bigg{|}<\epsilon$
@@
Hence, it suffices to choose $\alpha_{\epsilon}=\delta_{\epsilon}/d$, which implies
@@
$V(s,\mathbf{R_{acc}},k)\geq$ $\max_{a}\sum_{s^{\prime}}Pr(s^{\prime}|s,a)$
$\qquad V^{*}\bigg{(}s^{\prime},\mathbf{R_{acc}}+\gamma^{T-k}\mathbf{R}(s,a),k-1\bigg{)}-\epsilon-(k-1)\epsilon$
@@
A.4. Lemma 13 (Horizon Time)
Lemma 0.
Let $M$ be any MOMDP, and let $\pi$ be any policy in $M$. Assume the welfare function $W:\mathbb{R}^{d}\to\mathbb{R}$ is uniformly continuous and $\mathbf{R}(s,a)\in[0,1]^{d}\leavevmode\nobreak\ \forall s,a$. Then for all $\epsilon>0$, there exists $\delta_{\epsilon}>0$ such that if
$T\geq\left(\frac{1}{1-\gamma}\right)\log\left(\frac{\sqrt{d}}{\delta_{\epsilon}(1-\gamma)}\right)$
then for any state $s,$
Proof.
[…]
$=$ $\bigg{\|}\sum_{k=T+1}^{\infty}\gamma^{k-1}\mathbf{R}(s_{k},a_{k})\bigg{\|}$
$\leq$ $\gamma^{T}\left(\frac{\sqrt{d}}{1-\gamma}\right)$
Solving
$\gamma^{T}\left(\frac{\sqrt{d}}{1-\gamma}\right)\leq\delta_{\epsilon}$
Appendix B Necessity of Conditioning on Remaining Timesteps $t$
Our opening example in Figure 1 mentioned the necessity of conditioning on accumulated reward to model the optimal policy. However, the formulation of the optimal value function in Definition 7 also conditions on the number of time-steps remaining in the finite-time horizon task. One may naturally see this as undesirable in terms of adding computational overhead. Particularly in the continuing task setting with $\gamma$ very close to 1, it is natural to ask whether this conditioning is necessary.
Lemma 0.
In MORL with a nonlinear scalarization function $W$, for any constant $\alpha>0$, there exists an MOMDP such that any policy $\pi$ that depends only on the current state and accumulated reward (and does not condition on the remaining timesteps $t$) achieves expected welfare $V^{\pi}\leq\alpha V^{\pi^{*}}$, where $V^{\pi^{*}}$ is the expected welfare achieved by an optimal policy $\pi^{*}$ that conditions on $t$.
Proof.
We construct an MOMDP where the optimal action at a decision point depends critically on the remaining timesteps $t$. We show that any policy not conditioning on $t$ cannot, in general, approximate the optimal expected welfare within any constant factor.
Consider an MDP with states $s_{0}$, $s_{1}$, and $s_{2}$, where the agent chooses between actions $a$ and $b$ at $s_{0}$ and before reaching terminal states $s1$ or $s2$ with no further actions or rewards after reaching them. Choosing action $a$ results in a reward $\mathbf{R}(s_{0},a)=(\epsilon,R)$ and transition to $s_{1}$. Choosing action $b$ results in a reward $\mathbf{R}(s_{0},b)=(2\epsilon,0)$ and transition to $s_{2}$. The accumulated reward prior to reaching $s_{0}$ is $\mathbf{R}_{\text{acc}}=(0,D)$, where $D>0$. The scalarization function is $W(x_{1},x_{2})=x_{1}+x_{2}^{2}$, and we assume a discount factor $\gamma\in(0,1)$.
The agent must choose between actions $a$ and $b$ at $s_{0}$ to maximize the expected welfare:
$V=W\left(\mathbf{R}_{\text{acc}}+\gamma^{T-t}\mathbf{R}(s_{0},a\text{ or }b)\right).$
For action $a$:
$\mathbf{R}^{(a)}_{\text{total}}$ $=\mathbf{R}_{\text{acc}}+\gamma^{T-t}\mathbf{R}(s_{0},a)=\left(\gamma^{T-t}\epsilon,D+\gamma^{T-t}R\right),$
$V^{(a)}$ $=W\left(\mathbf{R}^{(a)}_{\text{total}}\right)=\gamma^{T-t}\epsilon+\left(D+\gamma^{T-t}R\right)^{2}.$
For action $b$:
$\mathbf{R}^{(b)}_{\text{total}}$ $=\mathbf{R}_{\text{acc}}+\gamma^{T-t}\mathbf{R}(s_{0},b)=\left(2\gamma^{T-t}\epsilon,D\right),$
$V^{(b)}$ $=W\left(\mathbf{R}^{(b)}_{\text{total}}\right)=2\gamma^{T-t}\epsilon+D^{2}.$
The difference in expected welfare is given by:
$\Delta V=V^{(a)}-V^{(b)}=\gamma^{T-t}\left(-\epsilon+2DR+\gamma^{T-t}R^{2}\right).$
When $T-t\to 0$ (i.e., $t$ is close to $T$, near the end of the horizon):
$\gamma^{T-t}\to 1,\quad\text{and}\quad\Delta V\to-\epsilon+2DR+R^{2}.$
Since $R$ is large, $2DR+R^{2}$ dominates, and $\Delta V>0$, making action $a$ optimal.
When $T-t$ is large (i.e., $t$ is small, far from $T$):
$\gamma^{T-t}\ll 1,\quad\text{and}\quad\Delta V\to\gamma^{T-t}\left(-\epsilon+2DR\right).$
If we choose $D=\frac{\epsilon-\delta}{2R}$ for some small $\delta>0$, then:
$-\epsilon+2DR=-\epsilon+2\left(\frac{\epsilon-\delta}{2R}\right)R=-\delta<0.$
Thus, $\Delta V<0$, making action $b$ optimal.
Any policy $\pi$ that does not condition on $t$ must choose either action $a$ or $b$ at $s_{0}$ regardless of $t$. If $\pi$ always chooses $a$, it will be suboptimal for large $T-t$, when action $b$ is optimal. If $\pi$ always chooses $b$, it will be suboptimal for small $T-t$, when action $a$ is optimal.
Let $\pi^{*}$ be the optimal policy that conditions on $t$. When action $a$ is optimal, we have
$\frac{V^{\pi}}{V^{\pi^{*}}}=\frac{2\gamma^{T-t}\epsilon+D^{2}}{\gamma^{T-t}\epsilon+\left(D+\gamma^{T-t}R\right)^{2}}.$
As $R\to\infty$, $V^{\pi^{*}}\to\left(D+\gamma^{T-t}R\right)^{2}$, which grows quadratically with $R$, while $V^{\pi}$ remains bounded. Therefore, the ratio approaches zero:
$\lim_{R\to\infty}\frac{V^{\pi}}{V^{\pi^{*}}}=0.$
When action $b$ is optimal, the analysis is analogous, leading to a similar conclusion.
Thus, for any constant $\alpha>0$, we can choose parameters $\epsilon,R,D,\gamma,T$ such that the expected welfare ratio $\frac{V^{\pi}}{V^{\pi^{*}}}\leq\alpha$. ∎
Appendix C Ablation Studies and Additional Experiments
C.1. Ablation Studies with $\gamma<1$
@@
We conduct further studies on the empirical effects of the discretization factor $\alpha$. For $W_{\text{SPF}}$, we set $\lambda=1$ (the smoothing factor). We gradually increase the disctretization interval of accumulated reward $\alpha$ gradually from $0.4$ to $1.2$ for a reward discount $\gamma=0.999$ for Taxi and $\gamma=0.99$ for Scavenger Hunt with the same settings as the experiments. As shown in Figure 4(a) and Figure 4(b), we observe that the solution quality is often very high (better than baselines considered in the main body) even for discretization values that are substantially greater than those needed for theoretical guarantees (with $\epsilon=1$, we need $\alpha=2.5\times 10^{-3}$). Nonetheless, the results exhibit non-monotonic behavior with respect to $\alpha$, which suggests that factors such as the alignment of discretization granularity with the welfare function’s smoothness play a role, especially in this large $\alpha$ regime where empirical performance exceeds the worst-case theoretical guarantees.
@@
Figure 4. Ablation study on discretization factor $\alpha$.
@@
This set of ablation studies demonstrates that the RAVI algorithm is capable of finding high-quality policies for optimizing expected scalarized return despite using a larger discretization parameter than that needed for our theoretical analysis. The practical takeaway is that, while smaller $\alpha$ values provide finer discretization and better theoretical guarantees, they incur increased computation costs. Conversely, larger $\alpha$ values can coarsen the discretization but still result in competitive performance, as demonstrated empirically.
C.2. Ablation Studies with $\gamma=1$
We also conduct further ablation studies on environments with $\gamma=1$ using coarser alpha values. We select $\alpha\in\{1,1.5,2,2.5,3\}$ and run 10 random initialization on each environment for each value. Note that the smallest $\alpha$ possible in both Taxi and Scavenger is $\alpha=1$ (which are the results we reported in the main text). Our results can be found at Table 2 and Table 3. We observe that there is a noticeable performance drop across both environments with a larger $\alpha$. Moreover, due to the nature of having a sparse reward signal in the Taxi environment, we observe no performance when $\alpha\geq 2.0$. On Scavenger, for $RD_{\text{threshold=2}}$, we observe that the agent is under-performing when $\alpha\geq 2.5$. This finding is also to be expected since the resources are scarce (only 6 resources in a $15\times 15$ grid), and enemies are plenty (33% of the grid world).
Table 2. Ablation study on discretization factor $\alpha$ with $\gamma=1$ on Taxi.
Dimension Alpha $W_{\text{Nash}}$ $W_{\text{Egalitarian}}$ $W_{p=-10}$ $W_{p=0.001}$ $W_{p=0.9}$
$d=2$ $\alpha=1.0$ 7.555$\pm$0.502 3.000$\pm$0.000 5.279$\pm$0.000 7.404$\pm$0.448 9.628$\pm$0.349
$\alpha=1.5$ 7.079$\pm$0.318 1.000$\pm$0.000 3.198$\pm$0.000 6.481$\pm$0.000 8.877$\pm$0.528
$\alpha=2.0$ 0.000$\pm$0.000 0.000$\pm$0.000 0.000$\pm$0.000 0.000$\pm$0.000 0.000$\pm$0.000
$\alpha=2.5$ 0.000$\pm$0.000 0.000$\pm$0.000 0.000$\pm$0.000 0.000$\pm$0.000 0.000$\pm$0.000
$\alpha=3.0$ 0.000$\pm$0.000 0.000$\pm$0.000 0.000$\pm$0.000 0.000$\pm$0.000 0.000$\pm$0.000
$d=3$ $\alpha=1.0$ 4.996$\pm$0.297 2.000$\pm$0.000 3.115$\pm$0.000 3.307$\pm$0.000 6.250$\pm$0.322
$\alpha=1.5$ 4.902$\pm$0.273 1.000$\pm$0.000 1.116$\pm$0.000 2.080$\pm$0.000 6.037$\pm$0.322
$\alpha=2.0$ 0.000$\pm$0.000 0.000$\pm$0.000 0.000$\pm$0.000 0.000$\pm$0.000 0.000$\pm$0.000
$\alpha=2.5$ 0.000$\pm$0.000 0.000$\pm$0.000 0.000$\pm$0.000 0.000$\pm$0.000 0.000$\pm$0.000
$\alpha=3.0$ 0.000$\pm$0.000 0.000$\pm$0.000 0.000$\pm$0.000 0.000$\pm$0.000 0.000$\pm$0.000
$d=4$ $\alpha=1.0$ 2.191$\pm$0.147 1.700$\pm$0.483 1.029$\pm$0.000 2.145$\pm$0.129 3.369$\pm$0.173
$\alpha=1.5$ 2.044$\pm$0.157 1.000$\pm$0.000 0.000$\pm$0.000 2.044$\pm$0.157 3.247$\pm$0.183
$\alpha=2.0$ 0.000$\pm$0.000 0.000$\pm$0.000 0.000$\pm$0.000 0.000$\pm$0.000 0.000$\pm$0.000
$\alpha=2.5$ 0.000$\pm$0.000 0.000$\pm$0.000 0.000$\pm$0.000 0.000$\pm$0.000 0.000$\pm$0.000
$\alpha=3$ 0.000$\pm$0.000 0.000$\pm$0.000 0.000$\pm$0.000 0.000$\pm$0.000 0.000$\pm$0.000
$d=5$ $\alpha=1.0$ 2.308$\pm$0.185 1.700$\pm$0.483 1.023$\pm$0.000 2.000$\pm$0.102 3.289$\pm$0.147
$\alpha=1.5$ 2.034$\pm$0.122 1.000$\pm$0.000 0.000$\pm$0.000 1.149$\pm$0.000 3.097$\pm$0.139
$\alpha=2.0$ 0.000$\pm$0.000 0.000$\pm$0.000 0.000$\pm$0.000 0.000$\pm$0.000 0.000$\pm$0.000
$\alpha=2.5$ 0.000$\pm$0.000 0.000$\pm$0.000 0.000$\pm$0.000 0.000$\pm$0.000 0.000$\pm$0.000
$\alpha=3.0$ 0.000$\pm$0.000 0.000$\pm$0.000 0.000$\pm$0.000 0.000$\pm$0.000 0.000$\pm$0.000
Table 3. Ablation study on discretization factor $\alpha$ with $\gamma=1$ on Scavenger.
Dimension Alpha $CD_{\rho=0.4}$ $RD_{\text{threshold}=2}$
$d=2$ $\alpha=1.0$ 1.336$\pm$0.240 3.400$\pm$0.843
$\alpha=1.5$ 1.236$\pm$0.279 3.600$\pm$0.699
$\alpha=2.0$ 0.197$\pm$0.381 0.100$\pm$0.316
$\alpha=2.5$ 0.400$\pm$0.578 -1.889$\pm$3.480
$\alpha=3.0$ 0.132$\pm$0.295 -12.500$\pm$22.405
Table 4. Runtime comparisons (in real-time hours).
Environment Dimension Function RAVI Mixture LinScal WelfareQ Mixture-M
Taxi $d=2$ $W_{\text{Nash}}$ 0.03 0.76 $\pm$ 0.20 1.98 $\pm$ 0.28 4.82 $\pm$ 1.65 0.15
$W_{\text{Egalitarian}}$ 0.01 0.69 $\pm$ 0.22 2.00 $\pm$ 0.31 2.61 $\pm$ 0.50 0.15
$W_{p=-10}$ 0.01 0.70 $\pm$ 0.24 2.11 $\pm$ 0.09 4.72 $\pm$ 1.81 0.15
$W_{p=0.001}$ 0.02 0.65 $\pm$ 0.29 2.10 $\pm$ 0.09 5.10 $\pm$ 1.75 0.15
$W_{p=0.9}$ 0.08 0.64 $\pm$ 0.21 2.11 $\pm$ 0.09 4.96 $\pm$ 1.77 0.15
$d=3$ $W_{\text{Nash}}$ 0.30 0.60 $\pm$ 0.29 2.14 $\pm$ 0.39 4.92 $\pm$ 1.91 0.32
$W_{\text{Egalitarian}}$ 0.01 0.72 $\pm$ 0.21 2.09 $\pm$ 0.30 2.26 $\pm$ 0.58 0.32
$W_{p=-10}$ 0.02 0.64 $\pm$ 0.27 2.18 $\pm$ 0.06 5.44 $\pm$ 1.91 0.32
$W_{p=0.001}$ 0.02 0.75 $\pm$ 0.11 2.10 $\pm$ 0.26 5.73 $\pm$ 1.80 0.32
$W_{p=0.9}$ 2.10 0.74 $\pm$ 0.28 2.18 $\pm$ 0.11 5.47 $\pm$ 1.79 0.32
$d=4$ $W_{\text{Nash}}$ 0.65 0.82 $\pm$ 0.14 2.19 $\pm$ 0.13 5.44 $\pm$ 1.62 0.53
$W_{\text{Egalitarian}}$ 0.06 0.71 $\pm$ 0.33 2.18 $\pm$ 0.13 2.47 $\pm$ 0.64 0.53
$W_{p=-10}$ 0.02 0.69 $\pm$ 0.20 2.12 $\pm$ 0.31 5.66 $\pm$ 1.58 0.53
$W_{p=0.001}$ 0.23 0.72 $\pm$ 0.23 2.21 $\pm$ 0.11 5.11 $\pm$ 1.80 0.53
$W_{p=0.9}$ 36.72 0.73 $\pm$ 0.22 2.21 $\pm$ 0.07 6.15 $\pm$ 1.78 0.53
$d=5$ $W_{\text{Nash}}$ 1.46 0.70 $\pm$ 0.21 2.25 $\pm$ 0.10 4.93 $\pm$ 1.72 0.75
$W_{\text{Egalitarian}}$ 0.28 0.71 $\pm$ 0.18 2.23 $\pm$ 0.05 2.37 $\pm$ 0.51 0.75
$W_{p=-10}$ 0.04 0.82 $\pm$ 0.08 2.21 $\pm$ 0.09 5.40 $\pm$ 1.83 0.75
$W_{p=0.001}$ 0.26 0.68 $\pm$ 0.17 2.05 $\pm$ 0.41 5.45 $\pm$ 1.77 0.75
$W_{p=0.9}$ 118.63 0.67 $\pm$ 0.27 2.16 $\pm$ 0.30 4.75 $\pm$ 1.59 0.75
Scavenger $d=2$ $CD_{p=0.4}$ 0.11 0.30 $\pm$ 0.05 0.25 $\pm$ 0.06 0.34 $\pm$ 0.05 4.71
$RD_{\text{threshold=2}}$ 0.09 0.25 $\pm$ 0.06 0.30 $\pm$ 0.05 0.29 $\pm$ 0.06 4.39
C.3. Visualizations of RAVI and Baselines
Given that Table 1 in the main text solely contains results evaluated from algorithms trained until convergence, it misses other important information such as the rate of convergence and the learning process of some online algorithms such as WelfareQ. Thus, in this subsection, we provide our visualizations of the learning curve for these algorithms. The plot is created using mean over 10 random initializations with standard deviation as shaded regions. Note that we use horizontal lines for model-based approaches as they do not have an online learning phase. Our results can be found at Figure 5, Figure 6, Figure 7, Figure 8, Figure 9, and Figure 10.
$d=2$
$d=3$
$d=4$
$d=5$
Figure 5. Taxi, $W_{\text{Nash}}$
$d=2$
$d=3$
$d=4$
$d=5$
Figure 6. Taxi, $W_{\text{Egalitarian}}$
$d=2$
$d=3$
$d=4$
$d=5$
Figure 7. Taxi, $W_{p=-10}$
$d=2$
$d=3$
$d=4$
$d=5$
Figure 8. Taxi, $W_{p=0.001}$
$d=2$
$d=3$
$d=4$
$d=5$
Figure 9. Taxi, $W_{p=0.9}$
$CD_{\rho=0.4}$
$RD_{\text{threshold}=2}$
Figure 10. Scavenger
Appendix D Details on the RAEE Algorithm